OpenStudy (anonymous):

What is the surface area of a circle that has a diameter of 14 inches

OpenStudy (anonymous):

multiply 14 by 3.14

OpenStudy (anonymous):

surface_Area =πr^2 = 3.14 * (7inches) * (7inches) = 3.14 * 49 inches square so the answer is approximately 153.86 inches square

OpenStudy (anonymous):

thank you I see what I did wrong